A landscape designer is designing a garden with three features: a sundial, a birdbath, and a rosebush. The sundial is due east of the birdbath; the rosebush is due south of the birdbath. The sundial is three times as far from the rosebush as it is from the birdbath. The distance between the rosebush and birdbath is 13.5 feet.
Answer by stanbon(75887)

A landscape designer is designing a garden with three features: a sundial, a birdbath, and a rosebush. The sundial is due east of the birdbath; the rosebush is due south of the birdbath. The sundial is three times as far from the rosebush as it is from the birdbath. The distance between the rosebush and birdbath is 13.5 feet.
-----
Plot the arrangement.
Put the sundial at (0,0)
Put the birdbath at (x,0)
Put the rosebush at (x,-3x)
------
Note:: sqrt[(x-0)^2+(3x-0)^2] = 13.5
x^2 + 9x^2 = 13.5^2
10x^2 = 182.5
x^2 = 18.25
x = 4.272 ft.
------------------
Cheers,
Stan H.
----------------